# ECHO: Efficient Dataset Condensation by Higher-Order Distribution Alignment

### 26 Dec 2023 v1

@@ 4.2 Comparison to the SOTA Methods
Model DSA MTT DM ECHO
10 ConvNet-3 52.1 56.4 48.9 63.5
ResNet-10 32.9 34.5 42.3 56.7
DenseNet-121 34.5 41.5 39.0 54.6
50 ConvNet-3 60.6 65.9 63.0 69.9
ResNet-10 49.7 43.2 58.6 66.6
DenseNet-121 49.1 51.9 57.4 66.1
Training Cost Evaluation. It is crucial to take into account the resource and time costs associated with dataset condensation, as discussed in prior works (Sajedi et al. 2023; Zhang et al. 2023). Some methods incur significantly higher time costs compared to the time required for training the entire dataset, making them a sub-optimal trade-off between efficacy and efficiency. We evaluate the time and peak GPU memory cost during condensing process of various baselines and our method. As shown in Tab. 5, our method demonstrates significant efficiency compared to optimization-oriented methods such as DC(Zhao and Bilen 2021b), DSA (Zhao and Bilen 2021a), and MTT (Cazenavette et al. 2022). Similar to DM (Zhao and Bilen 2023), our method scales effectively across varying IPCs. However, IDM, being a DM-based method, exhibits higher time costs compared to both DM (Zhao and Bilen 2023) and ECHO. The results in Tab. 5 can be well-explained by those in Tab. 6, where our method is observed to be as lightweight as DM (Zhao and Bilen 2023) and represents the first theoretically guaranteed DM-based method to align an infinite number of moments of the representation distributions.
Table 5: Time and GPU memeory cost comparison of SOTA datasets condensation methods. Run Time: the time for a single iteration. GPU memory: the peak memory usage during condensing. Both run time and GPU memory is averaged over 100 iterations. All experiments are conducted on CIFAR-10 with a single NVIDIA-A100 GPU. “✗” denotes out-of-memory issue.
IPC DC DSA DM MTT IDM ECHO
Run 1 0.16±${}_{\pm 0.01}$ 0.22±${}_{\pm 0.02}$ 0.08±${}_{\pm 0.02}$ 0.36±${}_{\pm 0.23}$ 0.50±${}_{\pm 0.03}$ 0.09±${}_{\pm 0.01}$
Time 10 3.31±${}_{\pm 0.02}$ 4.47±${}_{\pm 0.12}$ 0.08±${}_{\pm 0.02}$ 0.40±${}_{\pm 0.20}$ 0.48±${}_{\pm 0.05}$ 0.08±${}_{\pm 0.01}$
(Sec) 50 15.74±${}_{\pm 0.10}$ 20.13±${}_{\pm 0.58}$ 0.08±${}_{\pm 0.02}$ 0.58±${}_{\pm 0.04}$ 0.11±${}_{\pm 0.02}$
GPU 1 3515 3513 3323 2711 3223 2911
Memory 10 3621 3639 3455 8049 3179 3081
(MB) 50 4527 4539 3605 4027 3587
Table 6: Comparison of various DC methods.
Methods Theory No Model No Bi-Level
Guaranteed Training Optimization
DC -
Optimization DSA -
Oriented IDC -
MTT -
DM
DM CAFE
Based IDM
ECHO
Figure 4: (a) Ablation of the RKHS kernel. (b) Ablation of the number of iterations per model (IPM), the “Time” denotes the minimal time cost to get the corresponding accuracy, the experiments are conducted on CIFAR-10 with 10 images per class.
4.3 Ablation Study
Efficacy of Each Component. We utilize CIFAR-10 (Krizhevsky, Hinton et al. 2009) and SVHN (Netzer et al. 2011) with IPC=10 and IPC=50 to illustrate the impact of each component in our method. The corresponding results are presented in Tab. 7, revealing that each of the three components of ECHO enhances the performance of the condensed examples.
Table 7: Ablation study on CIFAR-10 and SVHN. “w/o RKHS” denotes using the loss of DM (Zhao and Bilen 2023). “w/o Multiple IPM” denotes the number of iterations per model is set to 1.
CIFAR-10 SVHN
IPC 10 50 10 50
ECHO w/o Factor 55.0 66.0 74.6 84.4
ECHO w/o RKHS 59.6 67.2 79.5 87.2
ECHO w/o Multiple IPM 62.0 68.7 82.2 88.4
ECHO 63.5 69.9 83.3 89.0
Sensitivity of the RKHS Kernel.  Different kernel functions construct distinct Reproducing Kernel Hilbert Spaces (RKHS). To investigate their influence, we adopt two additional kernel functions in addition to the Gaussian kernel: the linear kernel and the polynomial kernel. Fig. 3(a) illustrates the test accuracy under different kernel functions. As observed, the choice of kernel function $\mathcal{K}$ has minimal impact on the test accuracy of the condensed dataset. This indicates that as long as the selected kernel function is valid, ECHO can effectively embed the representation distributions in the constructed RKHS, resulting in a robust method.
@@ 4.3 Ablation Study
Impact of the IPM.  We conduct experiments using various number of iterations per model, and the corresponding performance is depicted in Fig. 3(b). We adopt CIFAR-10 with IPC=10 to showcase the impact of IPM. In addition to the test accuracy of condensed examples, we also provide the training time required to achieve the reported accuracy. As shown, increasing the number of IPM may lead to improved performance of the condensed data, but it also increases the training time. Conversely, an excessively large IPM can compromise the generalization ability of the condensed examples.
5 Conclusion
@@
In conclusion, this paper introduces a novel Distribution-Matching (DM)-based method called Efficient Dataset Condensation by Higher-Order Distribution Alignment (ECHO) for dataset condensation. With a theoretical guarantee, our method embeds the representation distributions of real and synthetic examples in a reproducing kernel Hilbert space, minimizing the maximum mean discrepancy between them to align their distributions in both first- and higher-order moments. Extensive experiments show the effectiveness and efficiency of our method. Notably, the efficiency of our method enables its application to more realistic and larger datasets. ECHO first studies the alignment of higher-order moments of the representation distributions between real and condensed examples, and establishes a strong baseline in DM-based methods for dataset condensation, which we believe will be valuable to the research community.
@@ Acknowledgments
This work was partially supported by grants from the National Key Research and Development Plan (2020AAA0140001), Beijing Natural Science Foundation (19L2040), and Open Research Project of National Key Laboratory of Science and Technology on Space-Born Intelligent Information Processing (TJ-02-22-01).
Appendix A Pseudo-Code of ECHO
The pseudo-code of ECHO is presented in Algorithm 1.
Algorithm 1 Efficient Dataset Condensation by Higher-Order Distribution Alignment (ECHO)
@@ Appendix A Pseudo-Code of ECHO
8:        Compute the ECHO loss $\mathcal{L}_{\text{ECHO}}$ between the sampled mini-batches according to Eq. (9)
9:        Update $\mathcal{S}$ as: $\mathcal{S}=\mathcal{S}-\eta\nabla_{\mathcal{S}}\mathcal{L}_{\text{ECHO}}$
@@ Appendix B Related Works
B.1 Optimization-Oriented Methods
@@
Parameter Matching.  The methods of parameter matching in dataset condensation is first proposed by Zhao et al. (Zhao and Bilen 2021b), which has been extended to various following works (Zhao and Bilen 2021a; Cazenavette et al. 2022; Kim et al. 2022). The key idea of parameter matching is that: the parameter induced by real and synthetic datasets should be consistent with each other. In DC (Zhao and Bilen 2021b), the resultant gradients w.r.t. the model on the synthetic dataset are encouraged to be close to that on the real dataset. DSA (Zhao and Bilen 2021a) further proposes to add the differentiable siamese augmentation before the images are feed to the model, which improved the informativeness of the synthetic dataset. Recognizing that single-step gradients may accumulate errors, MTT (Cazenavette et al. 2022) introduces a multi-step parameter matching method. This approach involves iteratively updating the synthetic images to ensure that the model trained on them follows similar trajectories towards convergence. On the other hand, there are some works focusing on effective model updating during the condensing process. The original work DC (Zhao and Bilen 2021b) uses the synthetic dataset to update the network, which could cause the overfitting problem in the eraly stage of training. To tackle this, IDC (Kim et al. 2022) proposes to update the model by real dataset, in which way the overfitting problem can be alleviated due to the large size of real dataset. Moreover, IDC (Kim et al. 2022) also proposed the factor technique, which can enrich the information of synthetic dataset by factoring and up-sampling. Though effective, these parameter matching methods are computationally expensive, as thousands of differently initialized networks are required to update the synthetic dataset. To accelerate the condensing process, Zhang et al. (Zhang et al. 2023) propose model augmentation, which adds a Gaussian perturbation on the early-stage models to save the time and storage cost to condense dataset.
B.2 Distribution-Matching-based Methods
@@ Appendix C Derivation of MMD
$\text{MMD}^{2}(\bm{P}_{\mathcal{T}},\bm{P}_{\mathcal{S}})=\lVert\mu[\bm{P}_{\mathcal{T}}]-\mu[\bm{P}_{\mathcal{S}}]\rVert^{2}$
$=\lVert\mu[\bm{P}_{\mathcal{T}}]\rVert^{2}+\lVert\mu[\bm{P}_{\mathcal{S}}]\rVert^{2}-2\lVert\mu[\bm{P}_{\mathcal{T}}]\rVert\lVert\mu[\bm{P}_{\mathcal{T}}]\rVert$
$=E^{2}_{\mathcal{T}}[\mathcal{K}(g_{\theta}(\bm{x}),\cdot)]+E^{2}_{\mathcal{S}}[\mathcal{K}(g_{\theta}(\bm{s}),\cdot)]$
@@
$=E_{\mathcal{T},\mathcal{T}}[\langle\mathcal{K}(g_{\theta}(\bm{x}),\cdot),\mathcal{K}(g_{\theta}(\bm{x}),\cdot)\rangle]$
$+E_{\mathcal{S},\mathcal{S}}[\langle\mathcal{K}(g_{\theta}(\bm{s}),\cdot),\mathcal{K}(g_{\theta}(\bm{s}),\cdot)\rangle]$
$-2E_{\mathcal{T},\mathcal{S}}[\langle\mathcal{K}(g_{\theta}(\bm{x}),\cdot),\mathcal{K}(g_{\theta}(\bm{s}),\cdot)\rangle]$
$=E_{\mathcal{T},\mathcal{T}}[\mathcal{K}(g_{\theta}(\bm{x}),g_{\theta}(\bm{x}))]+E_{\mathcal{S},\mathcal{S}}[\mathcal{K}(g_{\theta}(\bm{s}),g_{\theta}(\bm{s}))]$
$-2E_{\mathcal{T},\mathcal{S}}[\mathcal{K}(g_{\theta}(\bm{x}),g_{\theta}(\bm{s}))]\,\,//{\color[rgb]{0,0,1}\text{Reproducing property of $\mathcal{K}$}}$
@@ E.2 Dataset-Condensation
  • MTT (Cazenavette et al. 2022): matching the training trajectories of the models trained on real and synthetic images.
@@ Appendix F More Ablation Studies
Compatibility with Previous DM-based Methods.  The proposed kernel-form loss $\mathcal{L}_{\text{ECHO}}$ can be readily combined with previous DM-based methods. We substitute the loss function of DM (Zhao and Bilen 2023) and IDM (Zhao et al. 2023a) with our $\mathcal{L}_{\text{ECHO}}$, as shown in Tab. 8, the $\mathcal{L}_{\text{ECHO}}$ can lead to great improvements in the performance of the examples condensed by previous DM-based methods.
Table 8: Compatibility of ECHO with previous DM-based methods. All experiments are conducted on CIFAR-10 (Krizhevsky, Hinton et al. 2009).
@@
w/o $\mathcal{L}_{\text{ECHO}}$ w/ $\mathcal{L}_{\text{ECHO}}$ w/o $\mathcal{L}_{\text{ECHO}}$ w/ $\mathcal{L}_{\text{ECHO}}$
@@
Continual Learning.  We utilize our method to store more efficiently selected training samples in memory to address the challenge of catastrophic forgetting in continual learning (Masarczyk and Tautkute 2020). Following previous works (Zhao and Bilen 2023; Zhao et al. 2023a), the model is trained from scratch using only the latest memory, making the continual learning performance highly dependent on the quality of memory construction. To assess our method, we compare it with Random Selection, DM (Zhao and Bilen 2023), and DSA (Zhao and Bilen 2021a). The experimental setup involves class-incremental learning on the CIFAR-100 dataset, with a progressively increasing memory budget of 20 images per class. We conduct 5 and 10-step learning, randomly and evenly splitting the 100 classes into 5 and 10 learning steps, corresponding to 20 and 10 classes per step, respectively, which is consistent with the settings in DM (Zhao and Bilen 2023). The 3-layer ConvNet (Sagun et al. 2017) is employed in this experiment. As illustrated in Fig. 5, ECHO consistently outperforms other baselines in both scenarios, highlighting the superior quality of the condensed examples generated by ECHO for continual learning.
Figure 5: Application in continual learning. (a) 5-step (b) 10-step continual learning on CIFAR-100.
Appendix G Visualization Results
Fig. 6$\sim$9 visualize the synthetic images condensed by our method on low-resolution datasets with and without factor technique. We also provide the visualization of initialized and synthetic ImageNet-10 with 10 images per class in Fig. 10.

### 25 Feb 2024 v3

@@ Introduction
Model DSA DM M3D
10 ConvNet-3 52.1 48.9 63.5
ResNet-10 32.9 42.3 56.7
DenseNet-121 34.5 39.0 54.6
50 ConvNet-3 60.6 63.0 69.9
ResNet-10 49.7 58.6 66.6
DenseNet-121 49.1 57.4 66.1
[…]
Ablation Study
@@
Impact of the Iteration per Model (IPM).  We conduct experiments using various number of iterations per model, and the corresponding performance is depicted in Fig. 5(a). We adopt CIFAR-10 with 10 images per class to showcase the impact of IPM. In addition to the test accuracy of condensed examples, we also provide the training time required to achieve the reported accuracy. As shown, increasing the number of IPM may lead to improved performance of the condensed data, but it also increases the training time. Conversely, an excessively large IPM can compromise the generalization ability of the condensed examples.
Impact of the Kernel Function.  Different kernel functions construct distinct Reproducing Kernel Hilbert Spaces (RKHS). To investigate their influence, we adopt two additional kernel functions in addition to the Gaussian kernel: the linear kernel and the polynomial kernel. Fig. 5(b) illustrates the test accuracy under different kernel functions with 10 images per class. As observed, the choice of $\mathcal{K}$ has minimal impact on the performance of the condensed dataset. This indicates that as long as the selected kernel function is valid, our M3D can effectively embed the distributions in the constructed RKHS, resulting in a robust method.
Conclusion
@@
In conclusion, this paper introduces a novel Distribution-Matching (DM)-based method called M3D for dataset condensation. With a theoretical guarantee, our method embeds the representation distributions of real and synthetic examples in a reproducing kernel Hilbert space, minimizing the maximum mean discrepancy between them to align their distributions in both first- and higher-order moments. Extensive experiments show the effectiveness and efficiency of our method. Notably, the efficiency of our method enables its application to more realistic and larger datasets. M3D first studies the alignment of higher-order moments of the representation distributions between real and synthetic examples, and establishes a strong baseline in DM-based methods for dataset condensation, which we believe will be valuable to the research community.
@@ Acknowledgments
This work was partially supported by grants from the National Key Research and Development Plan (2020AAA0140001), and the Beijing Natural Science Foundation (19L2040), and the Open Research Project of National Key Laboratory of Science and Technology on Space-Born Intelligent Information Processing (TJ-02-22-01).
Appendix A Pseudo-Code of M3D
The pseudo-code of M3D is presented in Algorithm 1.
Algorithm 1 M3D: Dataset Condensation by Minimizing Maximum Mean Discrepancy
@@ Appendix A Pseudo-Code of M3D
8:        Compute the M3D loss $\mathcal{L}_{\text{M3D}}$ between the sampled mini-batches according to Eq. (9)
9:        Update $\mathcal{S}$ as: $\mathcal{S}=\mathcal{S}-\eta\nabla_{\mathcal{S}}\mathcal{L}_{\text{M3D}}$
@@ Appendix B Related Works
Optimization-Oriented Methods
@@
Parameter Matching.  The methods of parameter matching in dataset condensation is first proposed by Zhao et al. (Zhao and Bilen 2021b), which has been extended to various following works (Zhao and Bilen 2021a; Cazenavette et al. 2022; Kim et al. 2022). The key idea of parameter matching is that: the parameter induced by real and synthetic datasets should be consistent with each other. In DC (Zhao and Bilen 2021b), the resultant gradients w.r.t. the model on the synthetic dataset are encouraged to be close to that on the real dataset. DSA (Zhao and Bilen 2021a) further proposes to add the differentiable siamese augmentation before the images are feed to the model, which improved the informativeness of the synthetic dataset. Recognizing that single-step gradients may accumulate errors, MTT (Cazenavette et al. 2022) introduces a multi-step parameter matching method. This approach involves iteratively updating the synthetic images to ensure that the model trained on them follows similar trajectories towards convergence. On the other hand, there are some works focusing on effective model updating during the condensing process. The original work DC (Zhao and Bilen 2021b) uses the synthetic dataset to update the network, which could cause the overfitting problem in the eraly stage of training. To tackle this, IDC (Kim et al. 2022) proposes to update the model by real dataset, in which way the overfitting problem can be alleviated due to the large size of real dataset. Moreover, IDC (Kim et al. 2022) also proposed the factor technique, which can enrich the information of synthetic dataset by factoring and up-sampling. Though effective, these parameter matching methods are computationally expensive, as thousands of differently initialized networks are required to update the synthetic dataset. To accelerate the condensing process, (Zhang et al. 2023) propose model augmentation, which adds a Gaussian perturbation on the early-stage models to save the time and storage cost to condense dataset.
Distribution-Matching-based Methods
@@ Appendix C Derivation of MMD
$\text{MMD}^{2}(\bm{P}_{\mathcal{T}},\bm{P}_{\mathcal{S}})=\lVert\mu[\bm{P}_{\mathcal{T}}]-\mu[\bm{P}_{\mathcal{S}}]\rVert^{2}=\lVert\mu[\bm{P}_{\mathcal{T}}]\rVert^{2}+\lVert\mu[\bm{P}_{\mathcal{S}}]\rVert^{2}-2\lVert\mu[\bm{P}_{\mathcal{T}}]\rVert\lVert\mu[\bm{P}_{\mathcal{T}}]\rVert$
@@
$=E_{\mathcal{T},\mathcal{T}}[\langle\mathcal{K}(g_{\theta}(\bm{x}),\cdot),\mathcal{K}(g_{\theta}(\bm{x}),\cdot)\rangle]+E_{\mathcal{S},\mathcal{S}}[\langle\mathcal{K}(g_{\theta}(\bm{s}),\cdot),\mathcal{K}(g_{\theta}(\bm{s}),\cdot)\rangle]-2E_{\mathcal{T},\mathcal{S}}[\langle\mathcal{K}(g_{\theta}(\bm{x}),\cdot),\mathcal{K}(g_{\theta}(\bm{s}),\cdot)\rangle]$
$=E_{\mathcal{T},\mathcal{T}}[\mathcal{K}(g_{\theta}(\bm{x}),g_{\theta}(\bm{x}))]+E_{\mathcal{S},\mathcal{S}}[\mathcal{K}(g_{\theta}(\bm{s}),g_{\theta}(\bm{s}))]-2E_{\mathcal{T},\mathcal{S}}[\mathcal{K}(g_{\theta}(\bm{x}),g_{\theta}(\bm{s}))]\,\,//{\color[rgb]{0,0,1}\text{Reproducing property of $\mathcal{K}$}}$
@@ Appendix F More Ablation Studies
Compatibility with Previous DM-based Methods.  The proposed kernel-form loss $\mathcal{L}_{\text{M3D}}$ can be readily combined with previous DM-based methods. We substitute the loss function of DM (Zhao and Bilen 2023) and IDM (Zhao et al. 2023a) with our $\mathcal{L}_{\text{M3D}}$, as shown in Tab. 5, the $\mathcal{L}_{\text{M3D}}$ can lead to great improvements in the performance of the examples condensed by previous DM-based methods.
@@
w/o $\mathcal{L}_{\text{M3D}}$ w/ $\mathcal{L}_{\text{M3D}}$ w/o $\mathcal{L}_{\text{M3D}}$ w/ $\mathcal{L}_{\text{M3D}}$
@@
Table 5: Compatibility of M3D with previous DM-based methods. All experiments are conducted on CIFAR-10 (Krizhevsky, Hinton et al. 2009).
Appendix G More Visualization Results
Fig. 7$\sim$10 visualize the synthetic images condensed by our method on low-resolution datasets with and without the factor technique. We also provide the visualization of initialized and synthetic ImageNet-10 with 10 images per class in Fig. 11.